\underline{P398}

ULTRASENSITIVE DUAL-BEAM ABSORPTION SPECTROSCOPY, BASED ON WAVELENGTH-MODULATION TECHNIQUE AND LOGARITHMIC NOISE CANCELING

Yu. KURITSYN and \underline{V. LIGUER}, Institute of Spectroscopy, Russian Academy of Science, 142092 Troitsk Moscow Region, Russia.

Various registration methods for ultrasensitive diode laser based absorption spectroscopy are discussed. The sensitivity of a single-beam wavelength modulation technique is limited by excess laser noises and wavelengthdependent effects in optical system. A dual-beam logarithmic detection allows to reduce excess laser noises but actually shot-noise level is not usually achieved and sensitivity is limited by low-frequency transmittance fluctuations and a flicker noise from atomizer. Simultaneous using a double-beam logarithmic detection and a wavelength-modulation technique allows to minimize a low-frequency noise contribution. Absorbance of 7×10^{−8}/(Hz)^{−1/2} in such a system is demonstrated. Applications to trace measurements of chlorine in glowdischarge atomizer and lithium in graphite tube atomizer are described.

Time required: 15 min Session in which paper is recomended for presentation: 11 Comment: fax